How many milliliters of 0.116 M HCl acid will be needed to neutralize 25.00 mL of 0.0840 M Ba(OH)2 solution?

The glucose level of a patient was determined to be 130.0 mg/dL. How many milligrams of glucose would a 10.0 mL blood sample from this patient have?

A laboratory analysis of aspirin determines the following mass percent composition:

C 60.00%

H 4.48%

O 35.53%

Find the empirical formula.

A glass contains 50.0 mL of water. How many molecules of water are in the glass, if density of the water is 0.9977 g/mL?

How many moles of hydrogen molecules are in 15.69 g of the hydrogen?

sig figs please Thanks .
Answer #1

Question 1:

The balanced reaction is

Ba(OH)2 + 2HCl -------------- BaCl2 + 2H2O

Number of moles of HCl = 2 * number of moles of Ba(OH)2 [as per balanced reaction]

V/1000 * 0.116 = 2 * 25.00/1000 * 0.0840

V = 36.21 ml

Hence the volume of HCl needed will be 36.21 ml

Question 2:

Conversion factor: 1 dL = 100 ml

Amount of glucose in 10.00 ml blood sample = 10.00 ml * 130.0 mg/dL * 1 dL/100 ml = 13.0 mg

Hence the mass of glucose in 10.00 ml blood sample will be equal to 13.0 mg

Question 3:

Let the molar mass of compound be 100 g/mol

Number of moles of C = Mass/Molar mass = 60/12 = 5

Number of moles of H = Mass/Molar mass = 4.48/1.008 = 4.45

Number of moles of O = Mass/Molar mass = 35.53/16 = 2.22

Hence the ratio will be 2.22:4.45:5, which in simplified term can be written as 1:2:2.25 or 4:8:9

Hence the empirical formula of the compound will be C9H8O4

Question 4:

Mass of water = Density * Volume = 0.9977 g/ml * 50.00 ml = 49.885 g

Number of moles of H2O = Mass/molar mass = 49.885/18.015 = 2.77 moles

Number of molecules = 6.023 * 10^(23) * 2.077 = 1.67 * 10^(24) molecules

Question 5:

Molar mass of H2 = 2.016 g/mol

Number of moles of Hydrogen molecules = Mass/molar mass = 15.69/2.016 = 7.783 moles

Note - Post any doubts/queries in comments section.